# A new model of spontaneous scalarization of black holes induced by curvature and matter

### v1

@@ I Introduction
Black holes (BH) in GR are characterized by only three quantities: mass, spin and electric charge. No-hair theorems reinforce this claim and exclude extra scalar charges for some models when gravity couples to matter in the minimal way [10, 11]. Their proofs extend to many models [12], so the metric is like GR and the scalar field is trivial. However, if one assumption of the no-hair theorems is violated, a black hole can carry scalar hair. Early attempts led to pathological, unstable solutions (e.g. the Bocharova-Bronnikov-Melnikov-Bekenstein (BBMB) solution [13, 14]), but later analyses found regular solutions when the coupling meets certain conditions. Identifying which couplings lead to stable, regular solutions that match observations remains essential.
@@
GR has been extraordinarily successful in weak-field regimes, but its predictions in the strong-field regime remain largely untested. This experimental gap, now being addressed by gravitational-wave observations comes alongside theoretical puzzles that motivate extending GR. For example, phenomena like galaxy rotation curves (attributed to dark matter) and the accelerated cosmic expansion (dark energy) are not explained by GR with known fields. In addition, we cannot reconcile GR with quantum mechanics. These issues have prompted modifications of Einstein’s theory, such as adding higher-curvature terms to the action or introducing new degrees of freedom. Such extensions can emerge from fundamental physics, such as some attempts to unify gravity with other interactions naturally. A broad range of modified gravity models has been developed (see these review papers [15, 16, 17, 18, 19]), and upcoming strong-field tests offer a way to distinguish these alternative theories from GR in regimes where deviations may become pronounced.
@@
One of the principal frameworks for modifying gravity is the class of scalar-tensor theories that introduce a dynamical scalar field alongside the tensor field of GR. The prototypical example, Brans-Dicke theory, was formulated to incorporate Mach’s principle by allowing the gravitational constant to vary [20]. Modern scalar-tensor models are constructed to reduce to GR in appropriate limits (ensuring consistency with solar-system experiments) while offering richer behavior in large-scale structure or in other extreme conditions. In particular, they can provide an effective explanation for cosmic acceleration without a cosmological constant, all while obeying current local gravity constraints. The most general form of these theories is Horndeski theories [21, 22, 23, 24, 25]. They are designed so that their field equations remain second-order, avoiding ghost instabilities. No-hair theorems were developed for some classes of scalar-tensor theories. Hawking proved it for Brans-Dicke theories [10], then [16] extend the proof to $f(R)$ theories. Hui and Nicolis established no-hair theorems for most interesting classes of Horndeski theories [26]. Certain scalar-tensor theories allow the black hole to carry a nontrivial scalar field configuration. In other words they admit "hairy" black holes solutions that evade no-hair theorems. As proved in these theorems scalar hair is forbidden in many scalar-tensor models unless specific conditions are met, as when the scalar is coupled to additional invariants like Gauss-Bonnet (GB) term [27]. The possibility of scalar hair on black holes marks a significant departure from standard GR and motivates further empirical tests in the strong gravity regime.
@@
Higher-curvature corrections involving the GB invariant (a particular combination of squared curvature terms) arise naturally in several fundamental contexts. Notably, a GB term appears in the low-energy effective action of heterotic string theory [28, 29, 30, 31, 32] and in higher-dimensional generalizations of GR (Lovelock gravity) [33, 34]. In four dimensions the GB term by itself is a topological invariant and does not affect the field equations, but if it is coupled to a scalar field one obtains a nontrivial extension known as Einstein-scalar-Gauss-Bonnet (EsGB) gravity. EsGB models have second-order field equations and remain free of ghost instabilities, forming a subclass of Horndeski theories [23]. A well-studied example is Einstein-dilaton-Gauss-Bonnet (EdGB), which employs an exponential coupling $e^{-\alpha\phi}$ motivated by string theory . Hairy black hole solutions in this model were first found perturbatively in [35] and then numerically in [36]. After that [27] showed that shift symmetric coupling of the scalar field to GB invariant allow to get hairy solutions. Recently [37, 38] discovered a new violation of no-hair theorems in EsGB theories for a large class of coupling functions including a quadratic coupling . EsGB theories provide a promising arena to explore strong-gravity phenomena and to confront theoretical predictions with observational data.
@@
Another important category of alternative gravity involves a scalar field non-minimally coupled to the electromagnetic field. In such Einstein-Maxwell-Scalar (EMS) models, the scalar is coupled non-minimally to the Maxwell invariant through a coupling function $g(\phi)$. These models have very well motivated theoretical origins: Einstein-Maxwell-dilaton (EMD) theory for instance, emerges from Kaluza-Klein compactification and also appears as a low-energy limit of string theory, as well as in Supergravity frameworks. Another strong historical feature of such theories is that electromagnetic fields can source the scalar, leading to black hole solutions different from the usual Reissner-Nordström (RN) solution of GR. A well-known example is the Gibbons-Maeda-Garfinkle-Horowitz-Strominger dilaton black hole in EMD gravity [39, 40]. This theory does not admit the standard RN (bald) solution with scalar hair, any static charged BH must carry a "hair". The dilaton coupling thereby alters the black hole’s charge-mass relation; BH can attain charge-to-mass ratios exceeding the extremal limit of RN black holes (they are overcharged). Models with scalar-Maxwell couplings, often inspired by high energy theory, thus provide a rich framework for how additional fields might influence electromagnetic interactions in strong gravity, offering distinctive black holes signatures that can be probed observationally.
@@
Two broad classes of models realize this phenomenon: one is the curvature-induced scalarization through Gauss-Bonnet coupling [43, 44], the other is matter-induced scalarization through couplings to field such electromagnetism [45, 46]. In both cases, a nonminimal coupling of a scalar field $\phi$ to an additional invariant (geometric invariant such as the GB term, the Ricci scalar or the Chern-Simons term [47, 48, 49], or matter invariant such as the electromagnetic tensor $F_{\mu\nu}F^{\mu\nu}$) triggers a tachyonic instability when the invariant is sufficiently large, causing the GR solution (with $\phi=0$) to become unstable, inducing a new ground state with $\phi\neq 0$. This provides a natural way to circumvent no-hair theorems by introducing scalar hair only beyond a certain threshold of curvature or field strength. In the curvature-induced case, represented by EsGB theories, the scalar field is coupled to the Gauss–Bonnet invariant ${\cal G}$ (a quadratic curvature term). A representative action is
@@
with $f(\phi)$ a coupling function. The Schwarzschild black hole is a solution with $\phi=0$; however, if $f(\phi)$ is chosen such that $f^{\prime\prime}(0)\,{\cal G}>0$ induces a tachyonic mass for $\phi$, then above some coupling strength or below a certain mass scale the Schwarzschild solution becomes unstable. Indeed, it was shown that a static vacuum black hole can spontaneously scalarize when its curvature (characterized by $M^{-2}$ for a black hole of mass $M$) exceeds a critical value [44, 43]. The endpoint is a “hairy” black hole with a nontrivial $\phi(r)$ profile. A key feature of these Gauss–Bonnet models is that they preserve the GR solutions for weak curvature (large $M$), activating scalar hair only in the high-curvature regime. This behavior is motivated by the idea that deviations from GR appear only in extreme gravity environments (as might be expected from quantum gravity or high-energy effective corrections), thereby satisfying observational constraints in the weak-field regime, while providing new physics for compact objects. This is why curvature-induced scalarization of black holes has attracted a lot of interest [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Another attractive aspect is that the Gauss–Bonnet term is a natural curvature correction arising in effective descriptions (it appears in low-energy expansions of string theory), so EsGB models introduce scalar hair through a well-motivated extension of the gravitational action.
@@
Likewise, spontaneous scalarization can be induced by couplings to matter fields. An important example is the EMS class of models [45, 46, 69, 70], where the scalar field couples to the electromagnetic field invariant $F_{\mu\nu}F^{\mu\nu}$. In these models a charged Reissner–Nordström black hole (which in electrovacuum has $\phi=0$) can become unstable if the electric field is strong enough, leading to a new branch of scalarized charged black holes. For instance, with an action term of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$, the Maxwell field can trigger a tachyonic growth of $\phi$ when the charge $Q$ of the black hole exceeds a certain threshold. This matter-induced scalarization is analogous to the Gauss–Bonnet case: the scalar hair develops only when the electric field (or charge-to-mass ratio $Q/M$) is sufficiently large, ensuring that minimally coupled GR solutions remain stable for weakly charged black holes. Such EMS models have been extensively studied [71, 72, 73, 73, 74, 75, 61, 76, 77, 78, 79, 80, 81, 82, 83] as a way to generate hairy black holes that are smooth deformations of the Reissner–Nordström solution, again obeying the principle that hair appears only in regimes where Maxwell field is above some threshold.
@@
There have been some attempts to study charged black holes with Gauss-Bonnet-induced scalarization, without any direct scalar-Maxwell coupling [84, 85, 86]. These studies found that the inclusion of charge can enrich the scalarization behavior. For example, the Gauss–Bonnet term admits two types of scalarization (depending on the sign of the coupling) $GB^{+}$ or $GB^{-}$, and while a spherical symmetric neutral black hole only scalarizes in the positive-coupling case, adding charge allows scalarization with negative coupling ($GB^{-}$), but only near extremality. The electromagnetic field in those models influences the curvature-induced mechanism indirectly by modifying the black hole spacetime, but it does not act as an independent source of scalarization because $\phi$ does not couple to $F_{\mu\nu}F^{\mu\nu}$. Thus, the curvature-driven instability remains the primary route to hair in these models.
@@
A crucial challenge for the viability of EsGB scalarization is the stability of the scalarized solutions. It has become evident that many black holes with Gauss-Bonnet-induced hair suffer from perturbative instabilities [87]. In particular, for the simple choice of a quadratic coupling function $f(\phi)\propto\phi^{2}$, all the resulting scalarized black hole solutions are unstable under radial (spherically symmetric) perturbations [88, 89]. This means that although the hair can form, it does not represent a stable end state for evolution, small perturbations will drive the solution away (often back to the GR state). More sophisticated coupling functions (for example, exponential [43, 88] or higher-order forms [90, 89]) can produce branches of solutions that are radially stable. However, even those “stable” branches have been shown to be vulnerable to non-radial (angular) instabilities. In fact, modes with higher angular momentum can become unstable on hairy solutions that are stable to monopole perturbations. For instance, in theories of EsGB with Ricci coupled quadratically to $\phi$ refereed as (EsGBR) [91, 92], a Quadrupole ($l=2$) instability was found by [93], which can be extended to EsGB theories. Another instability was reported in EsGB and in EsGBR, called gradient instability, where the sound speed of angular propagation becomes imaginary for even perturbations in the case of large harmonic index $l$ [94]. The same authors later mentioned a mistake in their previous calculation, and found that the instability occurs for only part of the parameter space [95]. Many references reported a hyperbolicity loss in EsGB and EsGBR scalarization models and analyzed its threshold using numerical simulations [96, 97, 98, 99, 100, 101].
@@
Motivated by these developments, in this work we propose a new model of black hole spontaneous scalarization induced by both curvature and matter, aiming to capture both curvature-induced and matter-induced scalarization within a single framework. The model introduces a simultaneous coupling of the scalar field to the Gauss–Bonnet invariant and to $U(1)$ gauge field (a dark photon or an electromagnetic field), incorporating the two mechanisms within one framework. Our aim is to leverage the two sources of scalarization to overcome some of the difficulties faced by the pure EsGB case.
By including the matter-coupling directly, the scalar field has an additional channel to become tachyonic, which may widen the parameter space for stable scalarized solutions or provide new branches of hair that are absent in single-coupling models. Intuitively, the curvature coupling ensures that strong-gravity environments can trigger the scalar, while the matter coupling adds an extra contribution that may cure instabilities and modify the properties of the resulting black holes. This unified model is a logical next step because it reduces to the known cases in the appropriate limits (recovering pure Gauss-Bonnet scalarization when charge is zero, and pure Maxwell-induced scalarization when the Gauss-Bonnet coupling is turned off). At the same time, it offers the prospect of richer phenomenology by allowing a broader scalarization threshold for different masses and, potentially, more robust hairy black hole solutions. Another motivation, comes from the fact that this model can be seen as a subset of Maxwell-Horndeski theories, whose linear stability has already been studied [102]. This will allow us to check easily if the model suffers from the same instabilities reported previously. We should add that by analogy to similar cases this model might be seen as a low energy limit of string actions [35].
@@
In the following sections, we detail the construction of this model with two sources and explore its black hole solutions and their properties. Section II introduces the model and field equations, then analyzes the onset of scalarization, Section III presents numerical solutions of the scalarized black holes their physical properties like the scalar charge and thermodynamic properties, and Section IV discusses conclusions and possible future directions.
In this study, we adopt the convention $4\pi G=1=4\pi\epsilon_{0}$ for ease of calculation. The spacetime signature is (,+,+,+)(-,+,+,+)( - , + , + , + ). Our analysis is limited to spherically symmetric solutions, meaning that all metric and matter fields are functions of the radial coordinate only. To simplify our notation, any function initially introduced as dependent on the radial coordinate, like $X(r)$, will henceforth be referred to simply as $X$, keeping in mind its radial dependence. Derivatives are denoted as $X^{\prime}\equiv\frac{dX}{dr}$ for radial derivatives and $X_{,\phi}\equiv\frac{dX}{d\phi}$ for derivatives with respect to the scalar field $\phi$.

### v2

# New model of spontaneous scalarization of black holes induced by curvature and matter

@@ I Introduction
Black holes (BHs) in GR are characterized by only three quantities: mass, spin and electric charge. No-hair theorems reinforce this claim and exclude extra scalar charges for some models when gravity couples to matter in the minimal way [10, 11]. Their proofs extend to many models [12], so the metric is like GR and the scalar field is trivial. However, if one assumption of the no-hair theorems is violated, a black hole can carry scalar hair. Early attempts led to pathological, unstable solutions (e.g. the Bocharova-Bronnikov-Melnikov-Bekenstein (BBMB) solution [13, 14]), but later analyses found regular solutions when the coupling meets certain conditions. Identifying which couplings lead to stable, regular solutions that match observations remains essential.
@@
GR has been extraordinarily successful in weak-field regimes, but its predictions in the strong-field regime remain largely untested. This experimental gap, now being addressed by gravitational-wave observations comes alongside theoretical puzzles that motivate extending GR. For example, phenomena like galaxy rotation curves (attributed to dark matter) and the accelerated cosmic expansion (dark energy) are not explained by GR with known fields. In addition, we cannot reconcile GR with quantum mechanics. These issues have prompted modifications of Einstein’s theory, such as adding higher-curvature terms to the action or introducing new degrees of freedom. Such extensions can emerge from fundamental physics, such as some attempts to unify gravity with other interactions naturally. A broad range of modified gravity models has been developed (see the review papers [15, 16, 17, 18, 19]), and upcoming strong-field tests offer a way to distinguish these alternative theories from GR in regimes where deviations may become pronounced.
@@
One of the principal frameworks for modifying gravity is the class of scalar-tensor theories that introduce a dynamical scalar field alongside the tensor field of GR. The prototypical example, Brans-Dicke theory, was formulated to incorporate Mach’s principle by allowing the gravitational constant to vary [20]. Modern scalar-tensor models are constructed to reduce to GR in appropriate limits (ensuring consistency with solar-system experiments) while offering richer behavior in large-scale structure or in other extreme conditions. In particular, they can provide an effective explanation for cosmic acceleration without a cosmological constant, all while obeying current local gravity constraints. The most general form of these theories is Horndeski theories [21, 22, 23, 24, 25]. They are designed so that their field equations remain second-order, avoiding ghost instabilities. No-hair theorems were developed for some classes of scalar-tensor theories. Hawking proved it for Brans-Dicke theories [10], then [16] extended the proof to $f(R)$ theories. Hui and Nicolis established no-hair theorems for most interesting classes of Horndeski theories [26]. Certain scalar-tensor theories allow the black hole to carry a nontrivial scalar field configuration. In other words they admit "hairy" black holes solutions that evade no-hair theorems. As proved in these theorems scalar hair is forbidden in many scalar-tensor models unless specific conditions are met, such as when the scalar is coupled to additional invariants like the Gauss-Bonnet (GB) term [27]. The possibility of scalar hair on black holes marks a significant departure from standard GR and motivates further empirical tests in the strong gravity regime.
@@
Higher-curvature corrections involving the GB invariant (a particular combination of squared curvature terms) arise naturally in several fundamental contexts. Notably, a GB term appears in the low-energy effective action of heterotic string theory [28, 29, 30, 31, 32] and in higher-dimensional generalizations of GR (Lovelock gravity) [33, 34]. In four dimensions the GB term by itself is a topological invariant and does not affect the field equations, but if it is coupled to a scalar field one obtains a nontrivial extension known as Einstein-scalar-Gauss-Bonnet (EsGB) gravity. EsGB models have second-order field equations and remain free of ghost instabilities, forming a subclass of Horndeski theories [23]. A well-studied example is Einstein-dilaton-Gauss-Bonnet (EdGB), which employs an exponential coupling $e^{-\alpha\phi}$ motivated by string theory . Hairy black hole solutions in this model were first found perturbatively in [35] and then numerically in [36]. After that Ref [27] showed that shift symmetric coupling of the scalar field to GB invariant allow hairy solutions. Recently, Refs [37, 38] discovered a new violation of no-hair theorems in EsGB theories for a large class of coupling functions including a quadratic coupling. EsGB theories provide a promising arena to explore strong-gravity phenomena and to confront theoretical predictions with observational data.
@@
Another important category of alternative gravity involves a scalar field non-minimally coupled to the electromagnetic field. In such Einstein-Maxwell-Scalar (EMS) models, the scalar is coupled non-minimally to the Maxwell invariant through a coupling function $g(\phi)$. These models have very well motivated theoretical origins: Einstein-Maxwell-dilaton (EMD) theory for instance, emerges from Kaluza-Klein compactification and also appears as a low-energy limit of string theory, as well as in Supergravity frameworks. Another strong historical feature of such theories is that electromagnetic fields can source the scalar, leading to black hole solutions different from the usual Reissner-Nordström (RN) solution of GR. A well-known example is the Gibbons-Maeda-Garfinkle-Horowitz-Strominger dilaton black hole in EMD gravity [39, 40]. This theory does not admit the standard RN (bald) solution with scalar hair, any static charged BH must carry a "hair". The dilaton coupling thereby alters the black hole’s charge-mass relation; BH can attain charge-to-mass ratios exceeding the extremal limit of RN black holes (they are overcharged). Models with scalar-Maxwell couplings, often inspired by high energy theory, thus provide a rich framework for how additional fields might influence electromagnetic interactions in strong gravity, offering distinctive black hole signatures that can be probed observationally.
@@
Two broad classes of models realize this phenomenon: one is the curvature-induced scalarization through Gauss-Bonnet coupling [43, 44], and the other is matter-induced scalarization through couplings to the field such as electromagnetism [45, 46]. In both cases, a nonminimal coupling of a scalar field $\phi$ to an additional invariant (geometric invariant such as the GB term, the Ricci scalar or the Chern-Simons term [47, 48, 49], or matter invariant such as the electromagnetic tensor $F_{\mu\nu}F^{\mu\nu}$) triggers a tachyonic instability when the invariant is sufficiently large, causing the GR solution (with $\phi=0$) to become unstable, inducing a new ground state with $\phi\neq 0$. This provides a natural way to circumvent no-hair theorems by introducing scalar hair only beyond a certain threshold of curvature or field strength. In the curvature-induced case, represented by EsGB theories, the scalar field is coupled to the Gauss–Bonnet invariant $\cal G$} (a quadratic curvature term). A representative action is
@@
with $f(\phi)$ being a coupling function. The Schwarzschild black hole is a solution with $\phi=0$; however, if $f(\phi)$ is chosen such that $f^{\prime\prime}(0)\,{\cal G}>0$ induces a tachyonic mass for $\phi$, then above some coupling strength or below a certain mass scale the Schwarzschild solution becomes unstable. Indeed, it was shown that a static vacuum black hole can spontaneously scalarize when its curvature (characterized by $M^{-2}$ for a black hole of mass $M$) exceeds a critical value [44, 43]. The end point is a “hairy” black hole with a nontrivial $\phi(r)$ profile. A key feature of these Gauss–Bonnet models is that they preserve the GR solutions for weak curvature (large $M$), activating scalar hair only in the high-curvature regime. This behavior is motivated by the idea that deviations from GR appear only in extreme gravity environments (as might be expected from quantum gravity or high-energy effective corrections), thereby satisfying observational constraints in the weak-field regime, while providing new physics for compact objects. This is why curvature-induced scalarization of black holes has attracted a lot of interest [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Another attractive aspect is that the Gauss–Bonnet term is a natural curvature correction arising in effective descriptions (it appears in low-energy expansions of string theory), so EsGB models introduce scalar hair through a well-motivated extension of the gravitational action.
@@
Likewise, spontaneous scalarization can be induced by couplings to matter fields. An important example is the EMS class of models [45, 46, 69, 70], where the scalar field couples to the electromagnetic field invariant $F_{\mu\nu}F^{\mu\nu}$. In these models a charged Reissner–Nordström black hole (which in electrovacuum has $\phi=0$) can become unstable if the electric field is strong enough, leading to a new branch of scalarized charged black holes. For instance, with an action term of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$, the Maxwell field can trigger a tachyonic growth of $\phi$ when the charge $Q$ of the black hole exceeds a certain threshold. This matter-induced scalarization is analogous to the Gauss–Bonnet case: the scalar hair develops only when the electric field (or charge-to-mass ratio $Q/M$) is sufficiently large, ensuring that minimally coupled GR solutions remain stable for weakly charged black holes. Such EMS models have been extensively studied [71, 72, 73, 73, 74, 75, 61, 76, 77, 78, 79, 80, 81, 82, 83] as a way to generate hairy black holes that are smooth deformations of the Reissner–Nordström solution, again obeying the principle that hair appears only in regimes where the Maxwell field is above some threshold.
@@
There have been some studies of charged black holes with Gauss-Bonnet-induced scalarization, without any direct scalar-Maxwell coupling [84, 85, 86]. These studies found that the inclusion of charge can enrich the scalarization behavior. For example, the Gauss–Bonnet term admits two types of scalarization (depending on the sign of the coupling) $GB^{+}$ or $GB^{-}$, and while a spherical symmetric neutral black hole only scalarizes in the positive-coupling case, adding charge allows scalarization with negative coupling ($GB^{-}$), but only near extremality. The electromagnetic field in those models influences the curvature-induced mechanism indirectly by modifying the black hole spacetime, but it does not act as an independent source of scalarization because $\phi$ does not couple to $F_{\mu\nu}F^{\mu\nu}$. Thus, the curvature-driven instability remains the primary route to hair in these models.
@@
A crucial challenge for the viability of EsGB scalarization is the stability of the scalarized solutions. It has become evident that many black holes with Gauss-Bonnet-induced hair suffer from perturbative instabilities [87]. In particular, for the simple choice of a quadratic coupling function $f(\phi)\propto\phi^{2}$, all the resulting scalarized black hole solutions are unstable under radial (spherically symmetric) perturbations [88, 89]. This means that although the hair can form, it does not represent a stable end state for evolution, and small perturbations will drive the solution away (often back to the GR state). More sophisticated coupling functions (for example, exponential [43, 88] or higher-order forms [90, 89]) can produce branches of solutions that are radially stable. However, even those “stable” branches have been shown to be vulnerable to non-radial (angular) instabilities. In fact, modes with higher angular momentum can become unstable on hairy solutions that are stable to monopole perturbations. For instance, in theories of EsGB with Ricci coupled quadratically to $\phi$ referred to as (EsGBR) [91, 92], a quadrupole ($l=2$) instability was found by [93], which can be extended to EsGB theories. Another instability was reported in EsGB and in EsGBR, called gradient instability, where the sound speed of angular propagation becomes imaginary for even perturbations in the case of large harmonic index $l$ [94]. The same authors later mentioned a mistake in their previous calculation, and found that the instability occurs for only part of the parameter space [95]. Many references reported a hyperbolicity loss in EsGB and EsGBR scalarization models and analyzed its threshold using numerical simulations [96, 97, 98, 99, 100, 101].
@@
Motivated by these developments, in this work we propose a new model of black hole spontaneous scalarization induced by both curvature and matter, aiming to capture both curvature-induced and matter-induced scalarization within a single framework. The model introduces a simultaneous nonminimal coupling of the scalar field to the Gauss–Bonnet invariant and to $U(1)$ gauge field (a dark photon or an electromagnetic field), incorporating the two mechanisms within one framework. One should have in mind that a theory which contains a nonminimal coupling to both Gauss-Bonnet and Maxwell terms is very well motivated. Earlier works have consider it as a low energy limit of string actions [32, 28, 30, 35], and black hole solutions were already investigated in this effective string theory [35, 102, 103, 104, 105].
Our aim is to leverage the two sources of scalarization to try to overcome some of the difficulties faced by the pure EsGB case like instabilities. By including the matter-coupling directly, the scalar field has an additional channel to become tachyonic, which may widen the parameter space for stable scalarized solutions or provide new branches of hair that are absent in single-coupling models. Intuitively, the curvature coupling ensures that strong-gravity environments can trigger the scalar, while the matter coupling can also trigger scalarization and adds an extra contribution that modify the properties of the resulting black holes. This unified model is a logical next step because it reduces to the known cases in the appropriate limits (recovering pure Gauss-Bonnet scalarization when charge is zero, and pure Maxwell-induced scalarization when the Gauss-Bonnet coupling is turned off). At the same time, it offers the prospect of richer phenomenology by allowing a broader scalarization threshold for different masses and, potentially, more robust hairy black hole solutions.
Another motivation, comes from the fact that this model can be seen as a subset of Maxwell-Horndeski theories, whose linear stability has already been studied [106]. This will allow us to easily check if the model suffers from the same instabilities reported previously.
@@
In the following sections, we detail the construction of this model with two sources and explore its black hole solutions and their properties. Sec II introduces the model and field equations, then analyzes the onset of scalarization, Sec III presents numerical solutions of the scalarized black holes their physical properties like the scalar charge and thermodynamic properties, and Sec IV discusses conclusions and possible future directions.
In this study, we adopt the convention $4\pi G=1=4\pi\epsilon_{0}$ for convenience. The spacetime signature is (,+,+,+)(-,+,+,+). Our analysis is limited to spherically symmetric solutions, meaning that all metric and matter fields are functions of the radial coordinate only. To simplify our notation, any function initially introduced as dependent on the radial coordinate, like $X(r)$, will henceforth be referred to simply as $X$, keeping in mind its radial dependence. Derivatives are denoted as $X^{\prime}\equiv\frac{dX}{dr}$ for radial derivatives and $X_{,\phi}\equiv\frac{dX}{d\phi}$ for derivatives with respect to the scalar field $\phi$.